OpenStudy (anonymous):

the letter x and y represent rectangular coordinates, write the equation in polar coordinates (r,Ɵ): 2x²+2y²=3

OpenStudy (anonymous):

divide both sides by 2 and sub.s in r^2 = x^2 +y^2

OpenStudy (anonymous):

I have no idea O.O

OpenStudy (anonymous):

aight thanks for the help lol.

OpenStudy (anonymous):

i dont understand it !!

OpenStudy (anonymous):

lol its trig and pre calc

OpenStudy (anonymous):

Your welcome :DD haha sory im not the brightest crayon in the box in math XD

OpenStudy (anonymous):

oh well im in 7th grade soooo....

OpenStudy (anonymous):

hah.

OpenStudy (anonymous):

do you still have a question about this?

OpenStudy (anonymous):

what would the answer be? so if i got it right?

OpenStudy (anonymous):

r= sqrt(3/2)

OpenStudy (anonymous):

ok what about Ɵ?

OpenStudy (anonymous):

it's a circle... not a function of theta

OpenStudy (anonymous):

circles in polar coord.s are just a constant value of radius

OpenStudy (anonymous):

it's like a horizontal line in cartesian coord.s ... it doesn't matter what x is, y is always the same eg. y=2

OpenStudy (anonymous):

can you look at this and go to pg. 580 its # 67.